OpenStudy (data_lg2):

Consider the following system of equations 1. 4x+3y+3z=-8 2. 2x+y+z=-4 3. 3x-2y+(m^2-6)z)=m-4 Determine the value(s) of m for which this system of equations will have: a. no solution b. one solution c. an infinite number of solution I don't need answers, I just need some explanation..

OpenStudy (anonymous):

@SithsAndGiggles can you help me? please

OpenStudy (anonymous):

@phi @Hero @Luigi0210

OpenStudy (luigi0210):

You matrices

OpenStudy (luigi0210):

*use

OpenStudy (anonymous):

i'm not using matrices..

OpenStudy (phi):

It looks like sith is answering. This is not a trivial idea to explain... it depends on how much you already know.

OpenStudy (luigi0210):

Oops, nvm

OpenStudy (anonymous):

\[\begin{cases}4x+3y+3z=-8\\2x+y+z=-4\\3x-2y+(m^2-6)z=m-4\end{cases}\] \[\begin{pmatrix}4&3&3\\2&1&1\\3&-2&m^2-6\end{pmatrix}\begin{pmatrix}x\\y\\z\end{pmatrix}=\begin{pmatrix}-8\\-4\\m-4\end{pmatrix}\] I know that if the determinant of the coefficient matrix is non-zero, you have exactly one solution. Using a cofactor expansion (along the first column), you have \[\begin{align*}\begin{vmatrix}4&3&3\\2&1&1\\3&-2&m^2-6\end{vmatrix}&=4\begin{vmatrix}1&1\\-2&m^2-6\end{vmatrix}+2(-1)\begin{vmatrix}3&3\\-2&m^2-6\end{vmatrix}+3\begin{vmatrix}3&3\\1&1\end{vmatrix}\\ &=4\left(m^2-6-(-2)\right)-2\left(3m^2-18-(-6)\right)+3\left(3-3\right)\\ &=8-2m^2\end{align*}\] The determinant is zero when \(m=\pm2\), so one solution exists for \(m\in(-\infty,-2)\cup(-2,2)\cup(2,\infty)\). Plug in \(-2\) and \(2\) for \(m\) and see what happens to the system.

OpenStudy (anonymous):

@Data_LG2, do you mean you're currently not using matrices, or you're not supposed to?

OpenStudy (anonymous):

i'm not currently using matrices,..

OpenStudy (anonymous):

But you have learned about determinants and the like? It's best to use this method. Working this out by hand with elimination or substitution looks dreadful.

OpenStudy (anonymous):

I just did elementary operations but i'm stuck.. first, I eliminate X with equation 1 and 2, then I stopped when I'm calculating for eqn. 1 and 3..

OpenStudy (anonymous):

4x+3y+3z=-8 ...(1) 2x+y+z=-4 ....(2) \[3x-2y+\left( m ^{2} -6\right)z=m-4 ...(3)\] 2(2)-(1) gives, 4x+2y+2z-4x-3y-3z=-8+8 -y-z=0 y=-z 2(3)-3(2) gives, \[6x-4y+2\left( m ^{2}-6 \right)z-6x-3y-3z=2m-8+12\] \[-7y+\left( 2m ^{2}-12-3 \right)z=2m+4\] \[-7y+\left( 2m ^{2} -15\right)z=2m+4\] plug y=-z \[\left( 2m ^{2} -15+7\right)z=2m+4\] \[z=\frac{ 2\left( m+2 \right) }{ 2\left( m ^{2} -4\right) }\] z has no solution if \[m ^{2}-4=0,m ^{2}=4,m=\pm 2\] \[z=\frac{ m+2 }{ \left( m+2 \right)\left( m-2 \right) }=\frac{ 1 }{m-2 }\] it has infinite solutions by giving different value to m except -2 and 2